# Long-lived oscillations of false and true vacuum states in neutral atom systems

### v1

@@ IV Numerical Study
Figure 4: The Fourier and eigenvalue spectra for $\beta=\pm 0.02,0.06,0.10$. Left column: The Fourier magnitude $|\tilde{M}_{T}(\omega)|$ for several values of $\beta>0$, which yield an initial state close to the false vacuum. The curve opacities and vertical grey lines are described in Figure 3. The dashed lines show frequencies $\omega_{\ell v}>0$ and the dotted lines show frequencies $\omega_{\ell v}<0$. Middle column: The Fourier magnitude $|\tilde{M}_{T}(\omega)|$ for several values of $\beta<0$, which yield an initial state close to the true vacuum. Right column: The simultaneous eigenstates of $H$ and $T_{2}$. The conventions and color bar are described in Figure 3. In the Fourier spectra for $\beta>0$ and $\beta<0$, the secondary peaks shift in opposite directions, which can be used to distinguish between the false and true vacua, and understood from the trends in the eigenvalue spectra and our analysis in Section V.
IV.1 False vacuum case: $\beta>0$
We first consider an initial state close to the false vacuum state. We center our discussion around the case $\beta=0.01$, since it captures many important features and serves as an effective base case to describe trends with increasing $\beta$. Figure 3a shows the magnetization, which exhibits long-lived oscillations with several dominant frequencies. Figure 3b shows the Fourier spectrum of the transformed magnetization $M_{T}$ when it is evolved to different stop times $t_{\mathrm{Stop}}$. Naturally, a time-domain signal with a longer duration produces a frequency-domain signal with a higher frequency resolution. The stop times each increase by an order of magnitude to capture the improving resolution, and the lowest value $t_{\mathrm{Stop}}=4.0$ is closest to current experimental constraints for $\Omega/2\pi=1.0$ MHz (Section VI). The highest-resolution spectrum has a large, wide, dominant peak at $\omega\approx 19$ and a cluster of smaller, narrower, evenly-spaced secondary peaks at lower frequencies $\omega\in[11,16]$. The secondary peaks increase in amplitude with increasing frequency. Though not shown, the Néel OP has a similar Fourier spectrum to the magnetization.
@@ IV.1 False vacuum case: $\beta>0$
Figure 3c shows the simultaneous eigenstates of $H$ and $T_{2}$, with energy eigenvalues $E$ and translation eigenvalues $e^{ik}$ (equivalently, momenta $k$). We computed the eigenvalues by exact diagonalization with matrix balancing, a similarity transformation that improves the numerical stability [70, 71]. The figure shows only the low-energy 00- and $1$-meson eigenstates, those selected with energy $E/\Omega\leq 500$ and domain wall density $\rho n_{s}\lesssim 3$. The blue points show the false vacuum (FV) and true vacuum (TV) states; these have $k=0$ and $\rho n_{s}\lesssim 1$, i.e. approximately $\rho n_{s}\approx 0$ domain walls. We label the energy eigenvalues of false and true vacuum states as $E_{f}$ and $E_{t}$, respectively. The energy separation $\omega_{ft}=E_{f}-E_{t}$ between the two vacuum states is proportional to the local detuning parameter, $\omega_{ft}\propto\Delta_{\mathrm{loc}}$, as required since $\omega_{ft}\propto\mathcal{E}$. The red points show the 1-meson eigenstates; these have $1.5\lesssim\rho n_{s}\lesssim 2.5$, i.e. approximately $\rho n_{s}\approx 2$ domain walls.
@@
The band structure can be explained as follows. The parameters lie in the range $|\Delta_{\mathrm{loc}}|n_{s}\ll\Delta_{\mathrm{glob}}$ and $V_{2}\ll\Delta_{\mathrm{glob}}$, which has several consequences. The generation of two domain walls comes with an energy increase $\Delta_{\mathrm{glob}}$, which is much greater than the energy decrease possible from any domain size growth. The excited states thus split into separated manifolds, where the states in a single manifold approximately correspond to a single number $N$ of mesons. The Hamiltonian approximately conserves the meson number $N$. In the low-energy sector, manifolds with higher energy have larger $N$. In particular, the false and true vacua approximately correspond to 0-mesons, and the manifold of first excited states correspond to 1-mesons. The confinement field splits the manifold of 1-meson eigenstates into a tower of eight meson bands, approximately corresponding to the $n_{s}/2=8$ odd size domains $n=1,3,\ldots n_{s}-1$, as we discuss below in Section V. Each band splits into $n_{s}/2=8$ momenta $k$, since $T_{2}$ is a 2-site translation operator.
@@
In agreement with our model in Section III.2, we find that the peaks arise from interference between the 1-meson eigenstates with $k=0$ and the false vacuum state. In other words, the peaks are located at the frequencies $\omega_{\ell f}=E_{\ell}-E_{f}$, where $E_{\ell}$ for $\ell=1,\ldots,8$ are the energies of the $1$-meson eigenstates with momentum $k=0$, arranged in the order $E_{1}<\ldots<E_{8}$. The quench $\delta\beta=\beta-\beta_{\mathrm{prep}}$ conserves momentum, so it can only excite 1-meson eigenstates with the same momentum as the false vacuum, $k=0$. The quench can also excite a superposition of degenerate 1-meson eigenstates with momenta ±$\pm k$, but this is a higher order process. The observed frequencies support the 1-meson model in Equation 11.
@@
The frequencies $\omega_{\ell f}$ are shown by the vertical grey lines in Figure 3b. The dominant peak is located at the frequency $\omega_{8f}$. We can explain this as follows: out of all the 1-meson states $\left\lvert E_{\ell}\right\rangle$, the dominant basis elements in the state $\left\lvert E_{8}\right\rangle$ have the smallest Hamming distance from the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$ that dominates $\left\lvert E_{f}\right\rangle$, producing the largest matrix elements in Equation 13. The cluster of secondary peaks are located at the frequencies $\omega_{\ell f}$ for $\ell=1,\ldots,7$. They generally start with large amplitude at $\omega_{7f}$ and diminish in amplitude as they approach $\omega_{1f}$, eventually vanishing altogether. Again, we can explain this trend with the Hamming distance: as $\ell$ decreases from $7$ to $1$, the dominant basis elements in the states $\left\lvert E_{\ell}\right\rangle$ increase in Hamming distance from the $Z_{2}$ product state and overlap with the other basis states of $\left\lvert E_{f}\right\rangle$, which themselves decrease in amplitude with increasing Hamming distance from the $Z_{2}$ product state, so that the corresponding matrix elements decrease.
@@
We observe several trends in the spectrum as we increase $\beta>0$. These are captured in Figures 4a, 4d, 4g. The dominant peak shifts modestly to the left. The cluster of secondary peaks shift more rapidly to the left and spread out. The trends in the energy eigenvalue distribution provide a complementary picture, as shown in Figures 4c, 4f, 4i. The false and true vacuum energies separate, with the false vacuum energy moving upwards towards, and eventually through, the 1-meson energies. The 1-meson energies spread uniformly. The separation between $E_{f}$ and $E_{8}$ remains roughly constant, consistent with the dominant peak at $\omega_{8f}$ remaining at roughly the same location. The separation between $E_{f}$ and the remaining energies $E_{\ell}$ for $\ell=1,\ldots,7$ decreases, consistent with the secondary peaks at $\omega_{\ell f}$ shifting to the left.
IV.2 True vacuum case: $\beta<0$
@@
We next consider an initial state close to the true vacuum state, beginning with the case $\beta=-0.01$. Figure 3d shows the Fourier spectrum of $M_{T}$. The highest-resolution spectrum for $\beta=-0.01$ is similar to the one for $\beta=0.01$. However, the secondary peaks are located at higher frequencies $\omega\in[15,18]$ and they generally decrease in amplitude with increasing frequency. The simultaneous eigenstates of $H$ and $T_{2}$ for $\beta=-0.01$ are identical to those for $\beta=0.01$, since the Hamiltonian is invariant under the simultaneous transformation $\beta\rightarrow-\beta$ and $H\rightarrow T_{1}HT_{1}^{\dagger}$, where $T_{1}$ is the 1-site translation operator.
@@
The peaks for the case $\beta=-0.01$ are located at the frequencies $\omega_{\ell t}=E_{\ell}-E_{t}$, shown by the vertical grey lines in Figure 3d. The dominant peak is located at a frequency $\omega_{\sigma t}$ for a particular index $\sigma$, which can be explained as follows: out of all the states $\left\lvert E_{\ell}\right\rangle$, the dominant basis elements of the state $\left\lvert E_{\sigma}\right\rangle$ have the smallest Hamming distance from the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$ that dominates $\left\lvert E_{t}\right\rangle$, producing the largest matrix elements in Equation 13. The energy $E_{\sigma}$ can fall anywhere in the tower of 1-meson energy eigenvalues with $k=0$. The cluster of secondary peaks are located at the remaining frequencies $\omega_{\ell t}$ for $\ell\neq\sigma$. They generally start with large amplitude at $\omega_{1t}$ and diminish in amplitude as they approach $\omega_{8t}$ (excluding $\omega_{\sigma t}$), which can similarly be explained by the Hamming distance: as $\ell$ increases from $1$ to $8$ (excluding $\sigma$), the dominant basis elements in the states $\left\lvert E_{\ell}\right\rangle$ increase in Hamming distance from the $Z_{2}$ product state and overlap with the other basis states of $\left\lvert E_{t}\right\rangle$, which themselves decrease in amplitude with increasing Hamming distance from the $Z_{2}$ product state, so that the corresponding matrix elements decrease.
@@
The trends in the spectrum as we decrease $\beta<0$ are captured in Figures 4b, 4e, 4h. The dominant peak shifts modestly to the right. The cluster of secondary peaks shift more rapidly to the right and spread, eventually passing into and through the dominant peak. The trends in the energy eigenvalue distribution are consistent (Figures 4c, 4f, 4i). The true vacuum energy moves downwards. Though not explicitly labeled, the energy $E_{\sigma}$ transfers downward through the 1-meson energies. The separation between $E_{t}$ and $E_{\sigma}$ remains roughly constant, consistent with the dominant peak at $\omega_{8f}$ remaining at roughly the same location. The separation between $E_{t}$ and the remaining energies $E_{\ell}$ for $\ell\neq\sigma$ increases, consistent with the secondary peaks at $\omega_{\ell f}$ shifting to the right.
V Deconfinement From Rydberg Tails
@@
To explain the qualitative properties of the spectral peaks, though not the precise quantitative properties (i.e. locations, spacing, etc.), we examine the classical potentials of site basis states with 00- and $1$-meson excitations, as shown in Figure 5. We consider 1-meson states with $00$ domain walls only, since $11$ domain walls are suppressed by the Rydberg blockade. Let $n$ be the domain size; it must be odd given our restriction to $00$ domain walls. The next-nearest-neighbor interaction $V_{2}$ characterizes the leading effect of the Rydberg tails; we ignore higher order neighbor interactions, which are further suppressed. We note that our input parameter range implies $\Delta_{\mathrm{glob}}\gg V_{2}$ and $\Delta_{\mathrm{glob}}\gg|\Delta_{\mathrm{loc}}|$. In short, the qualitative properties of the spectral peaks arise due to competition between confinement from the local detuning $\Delta_{\mathrm{loc}}$ and deconfinement from the two-site Rydberg tails $V_{2}$.
@@
Figure 5: A schematic of the deconfinement model. The middle section shows several atom chains, where a red circle indicates a Rydberg state and a white circle indicates a ground state. The atoms are actually configured in circles as in Figure 2, but are depicted here in straight lines for convenience. The first chain shows the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$; it is a 0-meson state, which has no domain walls. The second, third, and fourth chains show 1-meson states, which have two $00$ domain walls, for domain sizes $n=1,3,5$. The Rydberg tails are dominated by the 2nd-nearest-neighbor interaction $V_{2}$. We apply a staggered local detuning field $\Delta_{\mathrm{loc},j}$, defined in Equation 3. The $Z_{2}$ product state is close to the false vacuum for $\Delta_{\mathrm{loc}}>0$; the corresponding pattern and results are shown in blue and in the left section. The $Z_{2}$ product state is close to the true vacuum for $\Delta_{\mathrm{loc}}<0$; the corresponding pattern and results are shown in teal and in the right section. The vertical arrows just left and right of the middle section show the order relations for the frequencies in each case, arising from those for the energies given in Equation 17 and Equations (18)-(19). The panels in the left and right sections show sample magnetization spectra for different relative values of $\Delta_{\mathrm{loc}}$ and $V_{2}$. The arrows below the panels indicate the qualitative relations between the frequencies and peaks. Our classical model of short range deconfinement thus shows that the dominant peak arises from the 1-meson state with $n=1$, the secondary peaks arise from the ordered 1-meson states with $n\geq 3$, the peaks shift in the direction opposite to the sign of $\Delta_{\mathrm{loc}}$, and the relative distribution of peaks is determined by competition between $\Delta_{\mathrm{loc}}$ and $V_{2}$.
@@
We first consider the case $\beta\propto\Delta_{\mathrm{loc}}>0$, shown in blue in Figure 5, i.e. the upper pattern in the middle section and the outcomes in the left section. The initial state is dominated by the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$, shown in the first chain in Figure 5, which is close to the false vacuum state. The $Z_{2}$ product state is a 0-meson state, i.e. it has no domain walls and thus domain size $n=0$. The classical potential $\epsilon_{0}\approx E_{f}$ of the 0-meson state is the zero point from which we will measure the 1-meson potentials, and is given by
@@
We next consider the case $\beta\propto\Delta_{\mathrm{loc}}<0$, shown in teal in Figure 5, i.e. the lower pattern in the middle section and the outcomes in the right section. The initial state is again dominated by the $Z_{2}$ product state, which is now close to the true vacuum state. The $Z_{2}$ product state has classical potential $\epsilon_{0}\approx E_{t}$. The expressions that result for the classical potentials and frequencies are the same as in the false vacuum case. However, since $\Delta_{\mathrm{loc}}<0$ now, they satisfy different order relations. The classical potentials satisfy the order relations
@@
The classical analysis applies similarly to the true vacuum case, producing some different results, captured in the panels on the right in Figure 5: ($2^{\prime}$) the secondary peaks arise from the frequencies $\omega_{[n\geq 3]}$, and these frequencies increase with increasing domain size $n$; ($3^{\prime}$) the secondary peaks decrease in amplitude with increasing frequency $\omega_{[n]}$ and thus increasing domain size $n$, since increasing the domain size produces states that are a greater Hamming distance $n$ from the $Z_{2}$ product state; and ($4^{\prime}$) as $\beta<0$ decreases, the peaks shift to higher frequencies and spread, since $\omega_{[n]}\sim-n\Delta_{\mathrm{loc}}$.

### v2

@@ IV Numerical Study
Figure 4: The Fourier and eigenvalue spectra for $\beta=\pm 0.02,0.06,0.10$. Left column: The Fourier magnitude $|\tilde{M}_{T}(\omega)|$ for several values of $\beta>0$, which yield an initial state close to the metastable state. The curve opacities and vertical grey lines are described in Figure 3. The dashed lines show frequencies $\omega_{\ell v}>0$ and the dotted lines show frequencies $\omega_{\ell v}<0$. Middle column: The Fourier magnitude $|\tilde{M}_{T}(\omega)|$ for several values of $\beta<0$, which yield an initial state close to the ground state. Right column: The simultaneous eigenstates of $H$ and $T_{2}$. The conventions and color bar are described in Figure 3. In the Fourier spectra for $\beta>0$ and $\beta<0$, the secondary peaks shift in opposite directions, which can be used to distinguish between the metastable and ground states, and understood from the trends in the eigenvalue spectra and our analysis in Section V.
IV.1 Metastable case: $\beta>0$
We first consider an initial state close to the metastable state. We center our discussion around the case $\beta=0.01$, since it captures many important features and serves as an effective base case to describe trends with increasing $\beta$. Figure 3a shows the magnetization, which exhibits long-lived oscillations with several dominant frequencies. Figure 3b shows the Fourier spectrum of the transformed magnetization $M_{T}$ when it is evolved to different stop times $t_{\mathrm{Stop}}$. Naturally, a time-domain signal with a longer duration produces a frequency-domain signal with a higher frequency resolution. The stop times each increase by an order of magnitude to capture the improving resolution, and the lowest value $t_{\mathrm{Stop}}=4.0$ is closest to current experimental constraints for $\Omega/2\pi=1.0$ MHz (Section VI). The highest-resolution spectrum has a large, wide, dominant peak at $\omega\approx 19$ and a cluster of smaller, narrower, evenly-spaced secondary peaks at lower frequencies $\omega\in[11,16]$. The secondary peaks increase in amplitude with increasing frequency. Though not shown, the Néel OP has a similar Fourier spectrum to the magnetization.
@@ IV.1 Metastable case: $\beta>0$
Figure 3c shows the simultaneous eigenstates of $H$ and $T_{2}$, with energy eigenvalues $E$ and translation eigenvalues $e^{ik}$ (equivalently, momenta $k$). We computed the eigenvalues by exact diagonalization with matrix balancing, a similarity transformation that improves the numerical stability [71, 72]. The figure shows only the low-energy 00- and $1$-meson eigenstates, those selected with energy $E/\Omega\leq 500$ and domain wall density $\rho n_{s}\lesssim 3$. The blue points show the metastable (MS) and ground (GS) states; these have $k=0$ and $\rho n_{s}\lesssim 1$, i.e. approximately $\rho n_{s}\approx 0$ domain walls. We label the energy eigenvalues of the metastable and ground states as $E_{m}$ and $E_{g}$, respectively. The energy separation $\omega_{mg}=E_{m}-E_{g}$ between the two states is proportional to the local detuning parameter, $\omega_{mg}\propto\Delta_{\mathrm{loc}}$, as required since $\omega_{mg}\propto\mathcal{E}$. The red points show the 1-meson eigenstates; these have $1.5\lesssim\rho n_{s}\lesssim 2.5$, i.e. approximately $\rho n_{s}\approx 2$ domain walls.
@@
The band structure can be explained as follows. The parameters lie in the range $|\Delta_{\mathrm{loc}}|n_{s}\ll\Delta_{\mathrm{glob}}$ and $V_{2}\ll\Delta_{\mathrm{glob}}$, which has several consequences. The generation of two domain walls comes with an energy increase $\Delta_{\mathrm{glob}}$, which is much greater than the energy decrease possible from any domain size growth. The excited states thus split into separated manifolds, where the states in a single manifold approximately correspond to a single number $N$ of mesons. The Hamiltonian approximately conserves the meson number $N$. In the low-energy sector, manifolds with higher energy have larger $N$. In particular, the metastable and ground states approximately correspond to 0-mesons, and the manifold of first excited states correspond to 1-mesons. The confinement field splits the manifold of 1-meson eigenstates into a tower of eight meson bands, approximately corresponding to the $n_{s}/2=8$ odd size domains $n=1,3,\ldots n_{s}-1$, as we discuss below in Section V. Each band splits into $n_{s}/2=8$ momenta $k$, since $T_{2}$ is a 2-site translation operator.
@@
In agreement with our model in Section III.2, we find that the peaks arise from interference between the 1-meson eigenstates with $k=0$ and the metastable state. In other words, the peaks are located at the frequencies $\omega_{\ell m}=E_{\ell}-E_{m}$, where $E_{\ell}$ for $\ell=1,\ldots,8$ are the energies of the $1$-meson eigenstates with momentum $k=0$, arranged in the order $E_{1}<\ldots<E_{8}$. The quench $\delta\beta=\beta-\beta_{\mathrm{prep}}$ conserves momentum, so it can only excite 1-meson eigenstates with the same momentum as the metastable state, $k=0$. The quench can also excite a superposition of degenerate 1-meson eigenstates with momenta ±$\pm k$, but this is a higher order process. The observed frequencies support the 1-meson model in Equation 11.
@@
The frequencies $\omega_{\ell m}$ are shown by the vertical grey lines in Figure 3b. The dominant peak is located at the frequency $\omega_{8m}$. We can explain this as follows: out of all the 1-meson states $\left\lvert E_{\ell}\right\rangle$, the dominant basis elements in the state $\left\lvert E_{8}\right\rangle$ have the smallest Hamming distance from the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$ that dominates $\left\lvert E_{m}\right\rangle$, producing the largest matrix elements in Equation 13. The cluster of secondary peaks are located at the frequencies $\omega_{\ell m}$ for $\ell=1,\ldots,7$. They generally start with large amplitude at $\omega_{7m}$ and diminish in amplitude as they approach $\omega_{1m}$, eventually vanishing altogether. Again, we can explain this trend with the Hamming distance: as $\ell$ decreases from $7$ to $1$, the dominant basis elements in the states $\left\lvert E_{\ell}\right\rangle$ increase in Hamming distance from the $Z_{2}$ product state and overlap with the other basis states of $\left\lvert E_{m}\right\rangle$, which themselves decrease in amplitude with increasing Hamming distance from the $Z_{2}$ product state, so that the corresponding matrix elements decrease.
@@
We observe several trends in the spectrum as we increase $\beta>0$. These are captured in Figures 4a, 4d, 4g. The dominant peak shifts modestly to the left. The cluster of secondary peaks shift more rapidly to the left and spread out. The trends in the energy eigenvalue distribution provide a complementary picture, as shown in Figures 4c, 4f, 4i. The metastable and ground state energies separate, with the metastable energy moving upwards towards, and eventually through, the 1-meson energies. The 1-meson energies spread uniformly. The separation between $E_{m}$ and $E_{8}$ remains roughly constant, consistent with the dominant peak at $\omega_{8m}$ remaining at roughly the same location. The separation between $E_{m}$ and the remaining energies $E_{\ell}$ for $\ell=1,\ldots,7$ decreases, consistent with the secondary peaks at $\omega_{\ell m}$ shifting to the left.
IV.2 Ground case: $\beta<0$
@@
We next consider an initial state close to the ground state, beginning with the case $\beta=-0.01$. Figure 3d shows the Fourier spectrum of $M_{T}$. The highest-resolution spectrum for $\beta=-0.01$ is similar to the one for $\beta=0.01$. However, the secondary peaks are located at higher frequencies $\omega\in[15,18]$ and they generally decrease in amplitude with increasing frequency. The simultaneous eigenstates of $H$ and $T_{2}$ for $\beta=-0.01$ are identical to those for $\beta=0.01$, since the Hamiltonian is invariant under the simultaneous transformation $\beta\rightarrow-\beta$ and $H\rightarrow T_{1}HT_{1}^{\dagger}$, where $T_{1}$ is the 1-site translation operator.
@@
The peaks for the case $\beta=-0.01$ are located at the frequencies $\omega_{\ell g}=E_{\ell}-E_{g}$, shown by the vertical grey lines in Figure 3d. The dominant peak is located at a frequency $\omega_{\sigma g}$ for a particular index $\sigma$, which can be explained as follows: out of all the states $\left\lvert E_{\ell}\right\rangle$, the dominant basis elements of the state $\left\lvert E_{\sigma}\right\rangle$ have the smallest Hamming distance from the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$ that dominates $\left\lvert E_{g}\right\rangle$, producing the largest matrix elements in Equation 13. The energy $E_{\sigma}$ can fall anywhere in the tower of 1-meson energy eigenvalues with $k=0$. The cluster of secondary peaks are located at the remaining frequencies $\omega_{\ell g}$ for $\ell\neq\sigma$. They generally start with large amplitude at $\omega_{1g}$ and diminish in amplitude as they approach $\omega_{8g}$ (excluding $\omega_{\sigma g}$), which can similarly be explained by the Hamming distance: as $\ell$ increases from $1$ to $8$ (excluding $\sigma$), the dominant basis elements in the states $\left\lvert E_{\ell}\right\rangle$ increase in Hamming distance from the $Z_{2}$ product state and overlap with the other basis states of $\left\lvert E_{g}\right\rangle$, which themselves decrease in amplitude with increasing Hamming distance from the $Z_{2}$ product state, so that the corresponding matrix elements decrease.
@@
The trends in the spectrum as we decrease $\beta<0$ are captured in Figures 4b, 4e, 4h. The dominant peak shifts modestly to the right. The cluster of secondary peaks shift more rapidly to the right and spread, eventually passing into and through the dominant peak. The trends in the energy eigenvalue distribution are consistent (Figures 4c, 4f, 4i). The ground state energy moves downwards. Though not explicitly labeled, the energy $E_{\sigma}$ transfers downward through the 1-meson energies. The separation between $E_{g}$ and $E_{\sigma}$ remains roughly constant, consistent with the dominant peak at $\omega_{8g}$ remaining at roughly the same location. The separation between $E_{g}$ and the remaining energies $E_{\ell}$ for $\ell\neq\sigma$ increases, consistent with the secondary peaks at $\omega_{\ell g}$ shifting to the right.
V Short-Range Meson Repulsion
@@
To explain the qualitative properties of the spectral peaks, though not the precise quantitative properties (i.e. locations, spacing, etc.), we examine the classical potentials of site basis states with 00- and $1$-meson excitations, as shown in Figure 5. We consider 1-meson states with $00$ domain walls only, since $11$ domain walls are suppressed by the Rydberg blockade. Let $n$ be the domain size; it must be odd given our restriction to $00$ domain walls. The next-nearest-neighbor interaction $V_{2}$ characterizes the leading effect of the Rydberg tails; we ignore higher order neighbor interactions, which are further suppressed. We note that our input parameter range implies $\Delta_{\mathrm{glob}}\gg V_{2}$ and $\Delta_{\mathrm{glob}}\gg|\Delta_{\mathrm{loc}}|$. In short, the qualitative properties of the spectral peaks arise due to competition between confinement from the local detuning $\Delta_{\mathrm{loc}}$ and short-range domain wall repulsion from the two-site Rydberg tails $V_{2}$.
@@
Figure 5: A schematic of the short-range domain wall repulsion model. The middle section shows several atom chains, where a red circle indicates a Rydberg state and a white circle indicates a ground state. The atoms are actually configured in circles as in Figure 2, but are depicted here in straight lines for convenience. The first chain shows the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$; it is a 0-meson state, which has no domain walls. The second, third, and fourth chains show 1-meson states, which have two $00$ domain walls, for domain sizes $n=1,3,5$. The Rydberg tails are dominated by the 2nd-nearest-neighbor interaction $V_{2}$. We apply a staggered local detuning field $\Delta_{\mathrm{loc},j}$, defined in Equation 3. The $Z_{2}$ product state is close to the metastable state for $\Delta_{\mathrm{loc}}>0$; the corresponding pattern and results are shown in blue and in the left section. The $Z_{2}$ product state is close to the ground state for $\Delta_{\mathrm{loc}}<0$; the corresponding pattern and results are shown in teal and in the right section. The vertical arrows just left and right of the middle section show the order relations for the frequencies in each case, arising from those for the energies given in Equation 17 and Equations (18)-(19). The panels in the left and right sections show sample magnetization spectra for different relative values of $\Delta_{\mathrm{loc}}$ and $V_{2}$. The arrows below the panels indicate the qualitative relations between the frequencies and peaks. Our classical model of short-range meson repulsion thus shows that the dominant peak arises from the 1-meson state with $n=1$, the secondary peaks arise from the ordered 1-meson states with $n\geq 3$, the peaks shift in the direction opposite to the sign of $\Delta_{\mathrm{loc}}$, and the relative distribution of peaks is determined by competition between $\Delta_{\mathrm{loc}}$ and $V_{2}$.
@@
We first consider the case $\beta\propto\Delta_{\mathrm{loc}}>0$, shown in blue in Figure 5, i.e. the upper pattern in the middle section and the outcomes in the left section. The initial state is dominated by the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$, shown in the first chain in Figure 5, which is close to the metastable state. The $Z_{2}$ product state is a 0-meson state, i.e. it has no domain walls and thus domain size $n=0$. The classical potential $\epsilon_{0}\approx E_{m}$ of the 0-meson state is the zero point from which we will measure the 1-meson potentials, and is given by
@@
We next consider the case $\beta\propto\Delta_{\mathrm{loc}}<0$, shown in teal in Figure 5, i.e. the lower pattern in the middle section and the outcomes in the right section. The initial state is again dominated by the $Z_{2}$ product state, which is now close to the ground state. The $Z_{2}$ product state has classical potential $\epsilon_{0}\approx E_{g}$. The expressions that result for the classical potentials and frequencies are the same as in the metastable case. However, since $\Delta_{\mathrm{loc}}<0$ now, they satisfy different order relations. The classical potentials satisfy the order relations
@@
The classical analysis applies similarly to the ground case, producing some different results, captured in the panels on the right in Figure 5: ($2^{\prime}$) the secondary peaks arise from the frequencies $\omega_{[n\geq 3]}$, and these frequencies increase with increasing domain size $n$; ($3^{\prime}$) the secondary peaks decrease in amplitude with increasing frequency $\omega_{[n]}$ and thus increasing domain size $n$, since increasing the domain size produces states that are a greater Hamming distance $n$ from the $Z_{2}$ product state; and ($4^{\prime}$) as $\beta<0$ decreases, the peaks shift to higher frequencies and spread, since $\omega_{[n]}\sim-n\Delta_{\mathrm{loc}}$.